### v2

# BioProBench: Comprehensive Dataset and Benchmark in Biological Protocol Understanding and Reasoning

Abstract
Biological protocols are fundamental to reproducibility and safety in life science research. While large language models (LLMs) perform well on general tasks, their systematic evaluation on these highly specialized, accuracy-critical, and inherently procedural texts remains limited. In this work, we present BioProBench, the first large-scale, multi-task benchmark for biological protocol understanding and reasoning. While there are several benchmark tasks involving protocol question answering, BioProBench provides a comprehensive suite of five core tasks: Protocol Question Answering, Step Ordering, Error Correction, Protocol Generation, and Protocol Reasoning, enabling a holistic evaluation of LLMs on procedural biological texts. Built upon 27K original protocols, it yields nearly 556K high-quality structured instances. We evaluate 12 mainstream open/closed-source LLMs. Experimental results reveal that some models perform well on basic understanding tasks (e.g., $\sim$70PQA-Acc., $>$64% ERR F1), but struggle significantly with deep reasoning and structured generation tasks like ordering and generation. Furthermore, model comparisons show diverse performance: certain open-source models approach closed-source levels on some tasks, yet bio-specific small models lag behind general LLMs, indicating limitations on complex procedural content. Overall, BioProBench, through its task design and experimental findings, systematically reveals the fundamental challenges for current LLMs in procedural knowledge understanding, deep adaptability to specific domains, reliability of structured reasoning, and handling of sophisticated precision and safety constraints, providing key directions for future AI in the field of scientific experiment automation. The code and data are available at: https://github.com/YuyangSunshine/bioprotocolbench and https://huggingface.co/datasets/BioProBench/BioProBench. footnotetext: Equal Contribution: $\star$; Corresponding author: $\dagger$.
1 Introduction
Biological experimental protocols, comprehensive documents detailing reagents, instruments, and crucially, step-by-step procedures, form the backbone of life science research, underpinning advances from drug discovery to synthetic biology. With laboratories increasingly adopting high-throughput automation murthy2024laboratory and cloud-based execution platforms boiko2023autonomous , the demand for reliable, automated interpretation and generation of these protocols has become increasingly critical ren2025towards . Even minor ambiguities or misinterpretations in protocols can lead to failed experiments, wasted resources, or, in the worst case, unsafe laboratory conditions.
Large language models (LLMs) have demonstrated significant progress across general biomedical text mining tasks nori2023can ; thirunavukarasu2023large ; lievin2024can ; yang2024advancing ; singhal2025toward , ranging from document classification and named-entity recognition to relation extraction, summarization, and question answering. While specialized models like BioBERT lee2020biobert , BioGPT luo2022biogpt , and BioMedGPT luo2023biomedgpt show domain-specific adaptation through fine-tuning on biomedical corpora, they typically lack the inherent architecture or training specifically geared towards handling complex procedural information like that found in protocols. Consequently, existing benchmarks for biomedical text processing, including BioASQ tsatsaronis2015overview , PubMedQA jin2019pubmedqa , LAB-Bench laurent2024lab and BixBench mitchener2025bixbench , despite driving progress in areas like question answering and data interpretation, have not systematically addressed the unique challenges of interpreting, reasoning about, and generating biological protocols, tasks central to experimental science.
(a) Data Sources and Counts
(b) Biology Category Counts
(c) New classes
(d) BioProBench Task and Sub-Task Counts
Figure 1: An overview of the BioProBench dataset and benchmark. (a) The data sources and corresponding raw data counts. (b) The biology categories of raw data counts. (c) The newly introduced classes in BioProBench. (d) The tasks and sub-tasks considered in BioProBench.
To fill this gap, we propose Biological Protocol Benchmark (BioProBench), the first large-scale, integrated multi-task dataset and benchmark for biological protocol understanding and reasoning, specifically designed for LLMs (Figure 1). BioProBench features a layered data design to support diverse model development stages: a raw corpus of 27K protocols for pretraining or RAG; a substantial downstream training set of over 550K structured instances across five fine-grained tasks for adaptation; and a held-out test set of 1,000 examples per task for benchmarking. This design provides both high-quality original corpora and focus on domain-specific fine-grained tasks. To accurately quantify procedural knowledge, we complement standard NLP metrics with novel domain-specific measures: keyword-based content metrics and embedding-based structural metrics. Evaluating 12 state-of-the-art LLMs (including open-source, closed-source, and domain-specific models) on BioProBench reveals fundamental limitations in their current ability to interpret, reason about and generate biological protocols. While advanced models like Gemini-2.5-pro-exp google2025gemini2_5_pro_exp achieves 70.27% PQA-ACC., and 65% ERR-F1, performance drops significantly on tasks requiring deeper procedural understanding (ORD-EM $\sim$ 50%) and structured generation(GEN-BLEU $<$ 15%). Furthermore, in ERR and GEN tasks, zero-shot and one-shot chain of thought (CoT) prompts are set to guide the correct reasoning process and improve performance. In addition, some open-source models approach closed-source levels on certain tasks, while smaller, domain-specific models often underperform due to insufficient capacity for capturing complex procedural dependencies. The fine-grained task settings and evaluation results provide important insights and directions for developing biomedical foundation models capable of professional-level experimental reasoning.
Overall, our contributions are summarized as follows:
  • We introduce BioProBench, the first large-scale, integrated multi-task dataset and benchmark for biological protocol understanding and reasoning, featuring over 556K high-quality structured instances derived from 27K high-quality protocols.
  • We design a comprehensive set of five tasks that challenge LLMs in different aspects of procedural comprehension and generation capabilities based on ambiguities, safety hazards, and nested dependencies commonly found in real-world biological experiments.
  • We propose a hybrid evaluation framework that combines standard NLP metrics with domain-specific metrics to accurately quantify model performance across all tasks.
  • Our extensive evaluation of 12 mainstream LLMs reveals significant limitations in their current ability to interpret, reason about, and generate biological protocols.
2 BioProBench Dataset and Benchmark Construction
2.1 Data Collection and Processing
As demonstrated in Figure. 1 (a),we collected 26,933 full-text protocols from 6 authoritative resources: Bio-protocol, Protocol Exchange, JOVE, Nature Protocols, Morimoto Lab, and Protocols.io. The details for these resources are provided in Appendix D. These protocols span 16 biological subfields. While Cell Biology and Imaging are predominant, the collection includes examples from Genomics, Immunology, Synthetic Biology, and more, ensuring broad domain coverage, as shown in Figures 1 (b) and (c). This broad distribution reflects the interdisciplinary nature of modern biological research and enhances the benchmark’s generalizability across diverse experimental contexts.
We performed initial processing, including deduplication and cleaning to remove formatting artifacts such as blank lines and HTML tags, using a combination of regular expressions and NLP techniques. Next, we performed structured extraction of key elements including protocol title, identifier (ID), keywords, and operation steps. To handle complex nested structures like sub-steps and nested lists, we applied parsing rules based on indentation and symbol levels. This process restored parent-child relationships, ensuring the accuracy of the extraction and laying a solid foundation for subsequent task generation. The overall workflow, including subsequent steps like task generation and quality control, is illustrated in Figure 2. Detailed data processing methods are provided in Appendix A.
Figure 2: Workflow of the BioProBench dataset construction. The process begins with collecting and structuring raw protocols and LLM-based data augmentation. This is followed by LLM-in-the-loop process to generate diverse task instances. Finally, tasks quality filtering through Self-Filtering mechanisms and Expert Feedback from domain specialists to ensure accuracy and relevance.
2.2 Task Instance Generation
Based on the enhanced data, we defined five core tasks: protocol question answering (QA), step ordering (ORD), error correction (ERR), protocol generation (GEN), and protocol reasoning (REA). These tasks are designed to provide fine-grained data for both training/fine-tuning and evaluation, addressing language ambiguities and operational risks inherent in real biological experiments, thereby facilitating LLMs’ development and robust benchmarking in this domain. Overall, BioProBench comprises 556,171 structured task instances. Their distribution and samples are illustrated in Figure 1 (d) and 3. Specifically, 380,697 instances originate from publicly licensed resources (Bio-protocol, Protocol Exchange, protocols.io) and will be released, while approximately 175,474 instances are derived (details descripted in Appendix D).
Figure 3: Samples across each tasks from BioProBench.
Protocol Question Answering (PQA) task simulates common information retrieval scenarios in biological experiments by querying three critical dimensions: reagent dosages, parameter values, and operational instructions. PQA addresses real-world ambiguities in protocol texts, including inconsistent units, undefined concentration ranges, and unordered procedural steps. Using Deepseek-V2 liu2024deepseek , we automatically construct multiple-choice questions with 5 options for each scenario segment and carefully designed the perturbation options to realistically reproduce the distractors that may be encountered in the laboratory. Three questions were generated for each dimension, introducing reasonable distractors and covering high-risk ambiguities. Detail format is in AppendiC.1.
Step Ordering(ORD) task aims to enhance the deep understanding of protocol hierarchy and procedural dependencies. Given the nested nature of biological protocols, we design two ordering forms: Top-Step (ordering main stages) and Child-Step (ordering sub-steps within a stage) to construct the dual challenges of the global and local levels. Instances are derived from original protocol steps and shuffled according to predefined rules (without using LLMs for shuffling) to create challenging ordering problems. The detail format is described in Appendix C.2.
Error Correction (ERR) task assesses the ability to identify and correct errors related to safety and result risks caused by text ambiguity or potential input mistakes. ERR focuses on typical deviations in reagents, parameters, and operations, simulating high-risk scenarios like volume overrides, incorrect concentrations, temperature conflicts, or reagent mislabeling. Using Deepseek-V3 liu2024deepseekv3 , we subtly modified key locations in original protocol steps to create error examples, generating an equal number of correct counterparts. Details on the format are in Appendix C.3.
Protocol Generation (GEN) task aims to enhance and evaluate the model’s ability to follow instructions under long text creation and professional constraints. To address challenges like incomplete descriptions and inconsistent terminology in real protocols, we extract key information from enhanced data and use Deepseek-R1 guo2025deepseek to automatically generate three tasks of different difficulty levels: Easy (atomic steps, no dependencies), Standard (two-level nesting, basic dependencies) and Difficult (three or more levels, complex dependencies). The examples generated by LLM must be strictly extracted from the original text rather than constructed. Up to five easy, three standard, and one difficult task are generated per protocol, retaining original step numbering. Priority is given to representative subtasks if source material is insufficient. Details on the format are in Appendix C.4.
Protocol Reasoning (REA) task introduces Chain of Thought (CoT) prompting based on ERR and GEN tasks to probe explicit reasoning pathways regarding experimental intent and potential risks. For GEN task, we designed a structured CoT template comprising <Objective>, <Precondition>, <Phase>, <Prameter>and <Structure>. This data guides the model to sort out the experimental objectives and conditional dependencies, ensuring that the generated content is both scientifically reasonable and in line with the field format specifications. For ERR task, CoT instances describe identified error types and analyze potential experimental risks and safety hazards. Details on the CoT format are in Appendix C.5.
2.3 Multi-stage quality control
To guarantee data reliability and quality, we implemente a multi-stage quality control process with a three-phase automated self-filtering pipeline. First, automated checks validate structural integrity (e.g. correct step indices, answer–question alignment). Second, template-based and semantic-similarity tests remove any mismatches between generated and reference answers. Third, outputs for challenging cases are regenerated and the best versions selected. Following automated filtering, a randomly selected 10% of instances underwent manual review by an expert group comprising over 5 PhD-level biologists. This manual review focuse on critical aspects such as experimental safety constraints, accuracy of technical terminology, and the scientific validity of procedural steps and reasoning chains. This multi-stage approach effectively balances dataset scale with the high fidelity required for evaluating complex biological protocol comprehension and generation.
3 Evaluation Metrics
We evaluate model performance using a combination of standard and domain-specific metrics. Table 1 provides an overview of all metrics used across different tasks. Details on standard metrics and LLM-as-a-Judge consistency are provided in the Appendix F. Below, we detail the novel domain-specific metrics we introduce: keyword-based content metrics and embedding-based structural metrics.
Keyword-Based Content Metrics.
The accurate inclusion of critical domain-specific terminology (e.g., reagents, equipment, specific techniques) is paramount in biological protocols. Standard overlap metrics might miss the significance of these key terms. Therefore, we employ a keyword-based analysis to specifically evaluate the model’s ability to capture the essential scientific concepts. We use KeyBERT keybert to extract the top $k=64$ keywords from reference ($K_{ref}$) and generated ($K_{gen}$) texts. We then compute Keyword Precision ($P_{K}=\frac{|K_{ref}\cap K_{gen}|}{|K_{gen}|}$), Keyword Recall ($R_{K}=\frac{|K_{ref}\cap K_{gen}|}{|K_{ref}|}$), and Keyword F1 ($F1_{K}=\frac{2\cdot P_{K}\cdot R_{K}}{P_{K}+R_{K}}$) based on the overlap between these keyword sets. This provides a targeted measure of whether the core scientific entities are correctly generated.
Table 1: Evaluation metrics for different tasks.
Task Standard Metrics Domain-specific Metrics
[…]
ORD Exact Match (EM) $\uparrow$, Kendall’s Tau ($\tau$) $\uparrow$, Failed $\downarrow$ -
ERR Precision (Prec.) $\uparrow$,Recall $\uparrow$, F1 $\uparrow$ -
[…]
REA Accuracy (Acc.) $\uparrow$, Precision (Prec.) $\uparrow$,Recall $\uparrow$, F1 $\uparrow$, Failed $\downarrow$ LLM-as-a-Judge Consistency (Consist.) $\uparrow$
Embedding-Based Structural Metrics.
Evaluating the structural fidelity of biological protocols generated by LLMs—specifically, ensuring necessary steps are present while minimizing extraneous ones—is crucial for usability but challenging for standard lexical overlap metrics. To address this, we utilize embedding-based metrics that operate at the step level. We represent the reference protocol as steps $S_{ref}=[s_{1},\dots,s_{n_{ref}}]$ and the generated protocol as $S_{gen}=[s^{\prime}_{1},\dots,s^{\prime}_{n_{gen}}]$. Each step is embedded using the “all-mpnet-base-v2” SentenceTransformer model  [^1] https://www.sbert.net/, and step similarity is measured using cosine similarity ($\text{Sim}(s,s^{\prime})$). We compute:
1) Step Recall (SR): This metric is motivated by the need to ensure completeness. It quantifies the proportion of essential reference steps that are semantically captured in the generated protocol, using a similarity threshold $\delta=0.7$.
$SR=\frac{\left|\{s\in S_{ref}:\max_{s^{\prime}\in S_{gen}}\text{Sim}(s,s^{\prime})\geq\delta\}\right|}{|S_{ref}|}$ (1)
2) Step Precision (SP): Complementary to SR, this metric addresses conciseness and relevance. It quantifies the proportion of generated steps that semantically correspond to a reference step, using a similarity threshold $\delta=0.7$. A higher SP indicates better quality (fewer spurious or irrelevant steps in the generation).
$SP=\frac{\left|{s^{\prime}\in S_{gen}:\max_{s\in S_{ref}}\text{Sim}(s,s^{\prime})\geq\delta}\right|}{|S_{gen}|}$ (2)
4 Experiments
We conducted experiments with a diverse range of LLMs to provide a broad perspective on current capabilities for biological protocol tasks. To this end, we selected 12 state-of-the-art models, encompassing three main categories: Closed-Source LLMs (o3-mini openai2025o3mini , GPT-4o achiam2023gpt , GPT-4-turbo openai2023gpt4turbo , Claude-3.7-sonnet anthropic2025claude37sonnet , Gemini-2.0-flash google2024gemini2 , Gemini-2.5-pro-exp google2025gemini2_5_pro_exp ), Open-Source LLMs (Deepseek-R1 guo2025deepseek , DeepSeek-V3 liu2024deepseek , QwQ-32B qwen2025qwq32b , Qwen-2.5-72b-instruct qwen2024qwen2_5_72b ), and Domain-Specific Bio-LLMs (BioMedGPT-10B luo2023biomedgpt , BioMistral-7B labrak2024biomistral , BioMistral-7B-DARE labrak2024biomistral ). For evaluation, we constructed a held-out test set consisting of approximately 1,000 representative samples per task. Full implementation details are provided in Appendix E.
4.1 Performance on Protocol Question Answering (PQA)
Table 2: Performance Comparison on PQA Task. The best value is highlighted in blue, and the runner-up value is highlighted in light blue.
Type Model Acc. $\uparrow$ BS $\downarrow$ Failed $\downarrow$
Closed source o3-mini openai2025o3mini 0.6567 0.2665 0.00%
GPT-4o achiam2023gpt 0.6350 0.2770 0.00%
GPT-4-turbo openai2023gpt4turbo 0.5792 0.3403 0.00%
Claude-3-7-sonnet anthropic2025claude37sonnet 0.6390 0.2454 0.50%
Gemini-2.0-flash google2024gemini2 0.6344 0.2605 0.17%
Gemini-2.5-pro-exp google2025gemini2_5_pro_exp 0.7027 0.1949 0.50%
Open source QwQ-32b qwen2025qwq32b 0.6367 0.2236 1.58%
Qwen2.5-72b-instruct qwen2024qwen2_5_72b 0.6530 0.2624 0.08%
Deepseek-V3 liu2024deepseek 0.6658 0.2172 0.00%
Deepseek-R1 guo2025deepseek 0.6783 0.1965 0.00%
Domain Specific BioMedGPT-10B luo2023biomedgpt 0.2021 0.5349 53.42%
BioMistral-7B labrak2024biomistral 0.2113 0.5705 33.75%
BioMistral-7B-DARE labrak2024biomistral 0.2794 0.5670 37.67%
On PQA task, evaluated LLMs show varied proficiency in comprehending and extracting information from biological protocols (Figure 4a and Table 2). The top performance is achieved by the closed-source model Gemini-2.5-pro-exp (70.27%Acc.). However, leading open-source models, notably Deepseek-R1 (67.83%Acc.) and Deepseek-V3 (66.58%Acc.), demonstrate strong capabilities, surpassing several other closed-source models. This highlights the increasing viability of advanced open-source models for domain-specific comprehension tasks, significantly narrowing the performance gap with closed-source alternatives in this domain. In contrast, the dedicated domain-specific models, BioMedGPT-10B and BioMistral-7B, performed exceptionally poorly (20-21%Acc.), exhibiting very high failure rates (33-53%Failed). This suggests that their limited parameter count and primary optimization for general biomedical text are insufficient for the broader procedural and technical understanding required by biological protocols, a capability currently better captured by larger, generalist LLMs. Accuracy generally correlated positively with confidence calibration, as measured by the Brier Score (BS). The top performer, Gemini-2.5-pro-exp, showed the best calibration (0.1949 BS), closely followed by Deepseek-R1 (0.1965 BS). Conversely, models with lower accuracy, such as GPT-4-turbo (57.92% Acc.), exhibited poorer calibration (0.3403 BS). This consistency suggests that models with stronger comprehension capabilities also provide more reliable confidence estimates for their predictions.
4.2 Performance on Step Ordering (ORD)
Table 3: Performance Comparison on ORD Task. The best value is highlighted in blue, and the runner-up value is highlighted in light blue.
[…]
Closed source o3-mini openai2025o3mini 0.4415 0.7333 2.07%
GPT-4o achiam2023gpt 0.4294 0.6268 7.32%
GPT-4-turbo openai2023gpt4turbo 0.3273 0.5278 23.43%
Claude-3-7-sonnet anthropic2025claude37sonnet 0.4781 0.7336 5.43%
Gemini-2.0-flash google2024gemini2 0.4021 0.6370 9.39%
Gemini-2.5-pro-exp google2025gemini2_5_pro_exp 0.5180 0.8104 4.22%
Open source QwQ-32b qwen2025qwq32b 0.4447 0.7047 5.68%
Qwen2.5-72b-instruct qwen2024qwen2_5_72b 0.4322 0.6572 7.92%
Deepseek-V3 liu2024deepseek 0.4005 0.6395 4.74%
Deepseek-R1 guo2025deepseek 0.4596 0.7453 2.93%
Domain Specific BioMedGPT-10B luo2023biomedgpt 0.0196 0.0203 91.21%
BioMistral-7B labrak2024biomistral 0.0417 0.0441 91.73%
BioMistral-7B-DARE labrak2024biomistral 0.0637 0.0486 64.86%
The ORD task directly probes the LLMs’ ability to reconstruct the logical sequence of biological protocols. This task proved highly challenging for all evaluated models, highlighting significant limitations in inferring complex procedural dependencies (Figure 4b and Table 3). Overall performance, measured by Exact Match (EM) accuracy requiring a perfectly ordered sequence, was low. The top-performing model, Gemini-2.5-pro-exp, achieved only 51.80% EM, indicating failure to capture the complete procedural flow in nearly half the cases. Other models lagged, with scores ranging down to 32.73% for GPT-4-turbo, reinforcing this bottleneck. However, Kendall’s Tau ($\tau$), measuring pairwise ranking correlation, presented a more positive view. $\tau$ scores were considerably higher, with Gemini-2.5-pro leading at 0.8104, and several others exceeding 0.73. This suggests models possess a reasonable grasp of local, pairwise step relationships even when failing to assemble the global sequence correctly.
4.3 Performance on Error Correction (ERR)
Table 4: Performance Comparison on ERR Task. The best value is highlighted in blue, and the runner-up value is highlighted in light blue.
[…]
Closed source o3-mini openai2025o3mini 0.6233 0.8443 0.2993 0.4420 0.00%
GPT-4o achiam2023gpt 0.6267 0.7500 0.3763 0.5011 0.00%
GPT-4-turbo openai2023gpt4turbo 0.5617 0.8000 0.1605 0.2674 0.00%
Claude-3-7-sonnet anthropic2025claude37sonnet 0.6093 0.7363 0.3367 0.4621 0.17%
Gemini-2.0-flash google2024gemini2 0.5867 0.7090 0.2893 0.4109 0.00%
Gemini-2.5-pro-exp google2025gemini2_5_pro_exp 0.6483 0.7009 0.5134 0.5927 0.00%
Open source QwQ-32b qwen2025qwq32b 0.6300 0.6552 0.5435 0.5941 0.00%
Qwen2.5-72b-instruct qwen2024qwen2_5_72b 0.5917 0.7500 0.2709 0.3980 0.00%
Deepseek-V3 liu2024deepseek 0.5858 0.7306 0.2676 0.3917 0.00%
Deepseek-R1 guo2025deepseek 0.6292 0.6197 0.6622 0.6403 0.00%
Domain Specific BioMedGPT-10B zhang2023biomedgpt 0.5150 0.5258 0.2726 0.3590 0.00%
BioMistral-7B labrak2024biomistral 0.4979 0.4978 0.5719 0.5323 1.42%
BioMistral-7B-DARE labrak2024biomistral 0.5008 0.4995 0.8311 0.6240 0.00%
The ERR task evaluates the ability to discern valid biological protocol steps from erroneous ones, a critical capability for ensuring experimental integrity. As detailed in Table. 4, models generally demonstrated varied success on this binary classification task, with overall accuracy hovering mostly between 58% and 65%. The closed-source Gemini-2.5-pro-exp achieved the highest accuracy (64.83%), indicating a reasonable ability to distinguish correct from incorrect steps overall. However, analyzing the error detection metrics reveals more nuanced performance and distinct strategies among models. A clear trade-off between Precision and Recall is evident across models. For instance, GPT-4-turbo exhibited high precision (80.00%) but very low recall (16.05%), suggesting it flags errors conservatively but is often correct when it does. Conversely, the open-source Deepseek-R1 achieved the highest recall among general-purpose models (66.22%), successfully identifying a majority of the errors, but with lower precision (61.97%). Notably, Deepseek-R1 achieved the highest F1 (64.03%), representing a strong balance. BioMedGPT-10B and BioMistral-7B showed poor performance, with accuracy near chance level ($\sim$50-51%) and modest F1. Interestingly, BioMistral-7B-DARE achieved exceptionally high Recall (83.11%), leading to a competitive F1 (62.40%), despite very low precision (49.95%).
4.4 Evaluating Reasoning Capabilities in Error Correction (REA-ERR)
Table 5: Performance Comparison on REA-ERR Task. The best value is highlighted in blue, and the runner-up value is highlighted in light blue. Consist. unreported for domain-specific models as structured outputs absent.
Type Model Acc. Prec. Recall F1 Failed Consist.
Closed source o3-mini openai2025o3mini 0.6505 0.8352 0.3729 0.5156 0.08% 0.2962
GPT-4o achiam2023gpt 0.6408 0.6803 0.5268 0.5938 0.00% 0.2945
GPT-4-turbo openai2023gpt4turbo 0.6033 0.7699 0.2910 0.4223 0.00% 0.1547
Claude-3-7-sonnet anthropic2025claude37sonnet 0.6508 0.7493 0.4498 0.5622 0.00% 0.2146
Gemini-2.0-flash google2024gemini2 0.6003 0.7542 0.2977 0.4269 0.33% 0.1780
Gemini-2.5-pro-exp google2025gemini2_5_pro_exp 0.6850 0.7273 0.5886 0.6506 0.00% 0.3943
Open source QwQ-32b qwen2025qwq32b 0.6421 0.5942 0.8943 0.7140 0.58% 0.2280
Qwen2.5-72b-instruct qwen2024qwen2_5_72b 0.6300 0.7175 0.4247 0.5336 0.00% 0.2529
Deepseek-V3 liu2024deepseek 0.6150 0.7906 0.3094 0.4447 0.00% 0.1980
Deepseek-R1 guo2025deepseek 0.6299 0.5903 0.8353 0.6917 0.25% 0.4526
Domain Specific BioMedGPT-10B luo2023biomedgpt 0.5194 0.5379 0.4000 0.4588 10.00% -
BioMistral-7B labrak2024biomistral 0.4735 0.4892 0.6276 0.5498 52.83% -
BioMistral-7B-DARE labrak2024biomistral 0.5124 0.5110 0.4964 0.5036 29.58% -
Table. 5 details models’ ability to identify errors and articulate their reasoning using one-shot CoT prompting. On this task, Gemini-2.5-pro-exp achieved the highest accuracy (68.50% Acc.), while QwQ-32b (71.40% F1) and Deepseek-R1 (69.17% F1) demonstrated strong balanced performance. Domain-specific models continued to exhibit significant limitations. Crucially, the Consist. metric, evaluating the validity of elicited reasoning (judged by Deepseek-V3), revealed Deepseek-R1 as the leader (45.26%), followed by Gemini-2.5-pro-exp (39.43%). This suggests Deepseek-R1’s structured reasoning more frequently produced justifications aligned with actual flaws.
A salient observation arises from comparing REA-ERR with the standard ERR task: structured reasoning significantly boosts the performance of QwQ-32b and Deepseek-R1. QwQ-32b’s F1 score increased markedly from 59.41% to 71.40%, making it the top F1 performer. Deepseek-R1 also achieved the second-highest F1 (69.17%) and the best Consist. score. These results suggest that these models, trained on complex reasoning tasks, can effectively utilize explicit guidance to improve error identification. However, a notable finding from the REA-ERR task is the large gap between Consist. and Acc. scores. This discrepancy highlights that, although structured reasoning improves overall performance, ensuring the coherence and validity of the reasoning process itself remains a crucial challenge for future research.
Table 6: Comprehensive Performance Comparison on GEN Task with direct and CoT Prompting. The best value is highlighted in blue, and runner-up value is highlighted in light blue.
Embedding Keywords N-gram
Type Model SR $\uparrow$ SP $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ BLEU $\uparrow$ METEOR $\uparrow$ ROUGE-L $\uparrow$ Failed $\downarrow$
Direct Prompting
Closed source o3-mini openai2025o3mini $0.3355$ $0.2539$ $0.1886$ $0.2999$ $0.2195$ $8.69$ $22.72$ $19.18$ $0.91$%
GPT-4o achiam2023gpt $0.3785$ $0.2492$ $0.1855$ $0.3344$ $0.2262$ $8.92$ $24.66$ $18.22$ $0.00$%
GPT-4-turbo openai2023gpt4turbo $0.3587$ $0.3172$ $0.1944$ $0.3475$ $0.2366$ $9.26$ $23.01$ $20.70$ $0.00$%
Claude-3-7-sonnet anthropic2025claude37sonnet $0.4280$ $0.2263$ $0.2167$ $0.3051$ $0.2392$ $8.38$ $24.78$ $16.83$ $0.00$%
Gemini-2.0-flash google2024gemini2 $0.3573$ $0.2440$ $0.1812$ $0.3413$ $0.2232$ $9.18$ $23.20$ $18.40$ $0.00$%
Gemini-2.5-pro-exp google2025gemini2_5_pro_exp $0.3704$ $0.2097$ $0.1716$ $0.3156$ $0.2105$ $7.11$ $24.48$ $14.85$ $0.00$%
Open source QwQ-32b qwen2025qwq32b $0.2979$ $0.2615$ $0.1902$ $0.3204$ $0.2262$ $8.40$ $23.05$ $18.56$ $0.00$%
Qwen2.5-72b-instruct qwen2024qwen2_5_72b $0.3723$ $0.2685$ $0.2045$ $0.3498$ $0.2457$ $10.23$ $23.97$ $19.91$ $0.00$%
Deepseek-V3 liu2024deepseek $0.3984$ $0.2584$ $0.1990$ $0.3097$ $0.2296$ $9.37$ $24.51$ $19.46$ $0.00$%
Deepseek-R1 guo2025deepseek $0.3491$ $0.2359$ $0.1421$ $0.2730$ $0.1758$ $8.62$ $23.85$ $18.33$ $0.00$%
Domain Specific BioMedGPT-10B luo2023biomedgpt $0.1941$ $0.1999$ $0.1721$ $0.2740$ $0.1967$ $6.84$ $21.18$ $15.15$ $0.00$%
BioMistral-7B labrak2024biomistral $0.0335$ $0.0445$ $0.1232$ $0.2170$ $0.1486$ $5.81$ $19.21$ $12.53$ $0.00$%
BioMistral-7B-DARE labrak2024biomistral $0.1400$ $0.2434$ $0.2163$ $0.1879$ $0.1756$ $4.66$ $15.12$ $15.90$ $0.00$%
CoT Prompting
Closed source o3-mini openai2025o3mini $0.3223$ $0.2572$ $0.1061$ $0.2040$ $0.1313$ $5.34$ $23.06$ $11.48$ $0.00$%
GPT-4o achiam2023gpt $0.3119$ $0.2437$ $0.1417$ $0.2643$ $0.1736$ $6.01$ $23.83$ $12.37$ $0.00$%
GPT-4-turbo openai2023gpt4turbo $0.3313$ $0.2591$ $0.1508$ $0.2855$ $0.1853$ $6.52$ $24.48$ $12.98$ $0.00$%
Claude-3-7-sonnet anthropic2025claude37sonnet $0.3918$ $0.2002$ $0.1408$ $0.2657$ $0.1732$ $5.72$ $24.30$ $11.56$ $0.00$%
Gemini-2.0-flash google2024gemini2 $0.3539$ $0.2489$ $0.1480$ $0.2807$ $0.1825$ $6.49$ $24.14$ $12.80$ $0.00$%
Gemini-2.5-pro-exp google2025gemini2_5_pro_exp $0.3579$ $0.2010$ $0.1406$ $0.2687$ $0.1736$ $4.79$ $22.81$ $10.12$ $0.00$%
Open source QwQ-32b qwen2025qwq32b $0.2745$ $0.2352$ $0.1542$ $0.2953$ $0.1906$ $6.25$ $23.46$ $13.31$ $0.00$%
Qwen2.5-72b-instruct qwen2024qwen2_5_72b $0.3377$ $0.2556$ $0.1509$ $0.2861$ $0.1858$ $6.89$ $24.91$ $13.67$ $0.00$%
Deepseek-V3 liu2024deepseek $0.3383$ $0.2767$ $0.1621$ $0.3094$ $0.2002$ $7.28$ $25.11$ $14.58$ $0.00$%
Deepseek-R1 guo2025deepseek $0.3208$ $0.2180$ $0.1450$ $0.2723$ $0.1780$ $5.00$ $21.37$ $11.04$ $0.00$%
4.5 Performance on Protocol Generation (GEN)
As Table 6 shows, N-gram metrics are uniformly low: BLEU tops out at 10.23 (Qwen2.5-72b-instruct) and METEOR at 24.78 (Claude-3-7-sonnet), while ROUGE-L at just 20.70 (GPT-4-turbo). These results indicate limited lexical overlap and fluency, with METEOR’s slightly higher score reflecting modest gains in synonym and paraphrase handling. Domain-specific metrics suggest a more pronounced performance gap. The highest Keyword F1 is only 0.2457, suggesting models grasp some core terms but lack consistency. SR remains below 0.43, meaning over half of the essential steps are omitted or misrendered. SP scores between 0.20 and 0.32 reveal frequent inclusion of irrelevant or extraneous steps, highlighting low precision in generated steps. Among models, Claude-3-7-sonnet leads on SR and METEOR, Qwen2.5-72b-instruct excels in BLEU and keyword Recall, and GPT-4-turbo performs best on ROUGE-L and SP.
Notably, employing CoT prompting consistently degraded performance across most models and metrics in the GEN task. Claude-3-7-sonnet’s SR dropped from 0.4280 to 0.3918, and its METEOR decreased slightly, while N-gram and Keyword scores saw more substantial reductions for most models under CoT. This suggests that untuned reasoning scaffolds can disrupt the generation of coherent, structured protocol text. To address this, BioProBench provides structured CoT exemplars, detailed in Section 2.2, that serve as a foundation for fine-tuning models to produce more fluent and logically consistent procedures. More experiments exploring the impact of CoT are in AppendixG.4.
5 Related Works
5.1 LLMs in Biomedical Domain Adaptation
The rapid advancement of large language models (LLMs)brown2020language ; achiam2023gpt ; anil2023palm ; team2023gemini ; touvron2023llama2 , fueled by Transformer architectures and self-supervised learning, has revolutionized natural language processing. Researchers in the biomedicine field face a critical challenge: the linguistic divide between general-purpose text corpora and the highly specialized lexicon of biomedical literaturejahan2024comprehensive . To mitigate this domain shift, three adaptation strategies have emerged: Architecture-specific tuning: BioBERT lee2020biobert adapts BERT through continued pre-training on PubMed resources, achieving 12.2-point MRR improvements in biomedical QA by capturing domain-specific syntax patterns. Task-oriented specialization: BioBART yuan2022biobart leverages BART’s encoder-decoder framework for generation-heavy tasks like literature summarization, while BioGPT luo2022biogpt employs GPT-style autoregressive modeling to attain state-of-the-art performance on complex reasoning benchmarks like PubMedQA jin2019pubmedqa . Multimodal integration: BioMedGPT luo2023biomedgpt pioneers cross-modal alignment between medical imaging and text, reducing radiology report errors to near-expert levels. Despite these advances, applying such models to experimental protocols reveals a persistent limitation. This is because technical details like reagent preparation sequences or equipment calibration steps, which are crucial in protocols, involve language operating under distinct constraints compared to clinical narratives or research abstracts, often eluding current models’ comprehension.
5.1.1 Biological Datasets and Benchmarks
The evaluation landscape for biomedical language models has evolved from basic QA (BioASQ tsatsaronis2015overview , PubMedQA jin2019pubmedqa ) to tasks probing more complex reasoning (DiseaseQA jin2021disease , ComprehensiveBioEval jahan2024comprehensive ). However, evaluations focusing specifically on biological protocols have been limited. For instance, LAB-Bench laurent2024lab includes many QA items but only 135 related to protocols. BixBench mitchener2025bixbench focuses on multi-step analysis over real-world datasets but does not specifically address protocol structure or procedural logic. In related clinical safety contexts, CARDBiomedBench bianchi2025cardbiomedbench assesses neurological diagnosis and risk mitigation, and MedXpertQA zuo2025medxpertqa ) assess domain-specific reasoning but not experimental protocols. While some emerging benchmarks address specific aspects of protocol processing, BioLP-bench ivanov2024biolp evaluates the diagnosis of injected errors in protocols, showing advanced models localize under 50% of errors. BioProt o2023bioplanner focuses on executable protocol generation through natural language to pseudocode mapping. Despite these advances, no current benchmark systematically evaluates the full spectrum of challenges inherent in real-world biological protocols, including strict sequencing, causal logic, precise quantitation, and safety compliance, across interpretation, reasoning, and generation tasks. BioProBench directly addresses this gap by providing a comprehensive, large-scale, multi-task benchmark focused specifically on end-to-end biological protocol understanding, reasoning, and generation.
6 Conclusion
Summary. In this work, we introduced BioProBench, the first large-scale, integrated multi-task benchmark for biological protocol understanding and reasoning. By curating 27K real-world protocols and generating over 556K structured task instances across five challenging tasks, alongside applying rigorous, domain-specific evaluation metrics, we exposed clear performance gaps in current large language models. Our evaluation exposed significant performance gaps in current LLMs, revealing struggles with deep temporal reasoning, causal logic, and generating safe, structured outputs, despite reasonable basic comprehension. By publicly releasing the dataset and evaluation suite, we offer a robust resource for advancing protocol-aware AI research.
Limitation & Future Work. Despite its comprehensive design, BioProBench has certain limitations. These include reliance on LLMs for task generation, potentially introducing subtle inaccuracies, and a focus solely on textual inputs, lacking real-world multimodal context (e.g., images, videos). Future work will address these by incorporating multimodal inputs and human-in-the-loop evaluation to improve reasoning depth, procedural accuracy, and practical applicability. We also aim to develop a protocol-specialized LLM by integrating retrieval-augmented generation (RAG) with structured chain-of-thought prompting, and applying lightweight domain adaptation techniques (e.g., LoRA) to efficiently adapt generalist models for handling complex procedural text.